Five years ago Roger was granted 5,000 incentive stock options and 3,500 nonqualified stock options. At that time the stock price was $20 per share.

The ISO exercise price was $20 per share and the NSO exercise price was $2 per share.

The value of the nonqualified options at the time of the grant was $18 per share. All of the ISOs and NSOs were exercised on the same day in the third year after the grant when the stock price was $42 per share. All of the option stock was then sold 14 months later for $65 per share.

  1. Will any part of the above transactions be taxed as ordinary income? If so, which?
  2. When Roger sold the stock, how were the proceeds taxed?
  3. Is the employer entitled to deduct any part of the above transactions? If so, which part(s) result in a deduction? How much is deductible?
  4. Calculate the amount of income from the grant to Roger.
Answer #1

Tax treatment of ISOs:

No tax implications of grant date and exercise date. Roger will recognize capital gain on the date of sale. Capital gain = (Sale price - purchase price) X number of shares = (65 - 20) X 5,000 = $225,000

No tax implications to employer for giving ISOs

Tax implications of non qualified stock options:

On grant date no tax implications.

On exercise date the difference between exercise price and the market price is treated as ordinary income to the employee and employer receives tax deduction for it.

Ordinary income = (Market price - Exercise price ) X number of shares = (42 - 2) X 3,5000 = $140,000

Employer deduction is also at $140,000

On sale excess of selling price over exercise price and ordinary income recognized on grant date is treated as capital gain.

Capital gain on sale = ( selling price - exercise price - ordinary income) X number of shares = (65 - 2 - 40) X 3,500 = 80,500

Will any part of the above transactions be taxed as ordinary income? If so, which?

Ordinary income recognized with respect to NSOs is $140,000

When Roger sold the stock, how were the proceeds taxed?

$225,000 long-term capital gain on sale of ISOs and 80,500 Long-term capital gain on sale of NSOs

Is the employer entitled to deduct any part of the above transactions? If so, which part(s) result in a deduction? How much is deductible?

$140,000 on exercise of NSOs

Calculate the amount of income from the grant to Roger.

Income on grant to roger is = Sale price - purchase price = 5,000 X 65 + 3,500 X 65 - 5,000 X 20 - 3,500 X 2 = 552,500 - 107,000 = 445,500